# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ III Preliminaries
Definition III.3 (PBA)
[…]
  • $\mathcal{Q}_{\mathcal{P}}={\mathcal{S}}_{r}\times\mathcal{Q}_{{\mathcal{B}}}$ is the set of product states;
  • $\mathcal{Q}_{\mathcal{P}}^{0}=\{s_{r}^{0}\}\times\mathcal{Q}_{\mathcal{B}}^{0}$ is a set of initial states;
  • $\longrightarrow_{{\mathcal{P}}}\subseteq\mathcal{Q}_{\mathcal{P}}\times\mathcal{Q}_{\mathcal{P}}$ is the transition relation defined by the rule: $\frac{(s_{r}\rightarrow_{r}s^{\prime}_{r})\wedge(q\xrightarrow{{\mathcal{L}}_{r}\left(s_{r}\right)}_{{\mathcal{B}}}q^{\prime})}{q_{{\mathcal{P}}}=\left(s_{r},q\right)\longrightarrow_{\mathcal{P}}q_{{\mathcal{P}}}^{\prime}=\left(s^{\prime}_{r},q^{\prime}\right)}$. The transition from the state $q_{\mathcal{P}}\in\mathcal{Q}_{\mathcal{P}}$ to $q_{\mathcal{P}}^{\prime}\in\mathcal{Q}_{\mathcal{P}}$, is denoted by $(q_{\mathcal{P}},q_{\mathcal{P}}^{\prime})\in\longrightarrow_{\mathcal{P}}$, or $q_{\mathcal{P}}\longrightarrow_{\mathcal{P}}q_{\mathcal{P}}^{\prime}$;
  • $\mathcal{Q}_{\mathcal{P}}^{F}={\mathcal{S}}_{r}\times\mathcal{Q}_{\mathcal{B}}^{F}$ is a set of accepting/final states;
IV Hierarchical Syntactically Co-safe LTL
In this work, we focus on a particular subset of LTL formulas known as syntactically co-safe LTL, or sc-LTL for short [65]. As established by [65], any LTL formula encompassing only the temporal operators $\Diamond$ and $\mathcal{U}$ and written in positive normal form (where negation is exclusively before atomic propositions) is classified under syntactically co-safe formulas. This category does not include the $\square$ operator. Sc-LTL formulas can be satisfied by finite sequences followed by any infinite repetitions. This characteristic makes sc-LTL apt for modeling and reasoning about systems with finite durations, such as those found in the robotics field. We first present the syntax of hierarchical sc-LTL, then we propose the interpretation of semantics of hierarchical sc-LTL specifications, along with the analysis of the expressiveness.
IV-A Syntax of Hierarchical sc-LTL
Definition IV.1 (Hierarchical sc-LTL)
Hierarchical sc-LTL is structured into $K$ levels, labeled as $L_{1},\ldots,L_{K}$, arranged from the highest to the lowest. Each level $L_{k}$, where $k\in[K]$, contains $n_{k}$ sc-LTL formulas. The hierarchical sc-LTL can be represented as $\Phi=\left\{\phi_{k}^{i}\,|\,k\in[K],i\in[n_{k}]\right\}$, where $\phi_{k}^{i}$ denotes the $i$-th sc-LTL formula at level $L_{k}$. The hierarchical sc-LTL adheres to the following rules:
  1. 1.
    Each formula at a given level $L_{k}$, for $k\in[K-1]$, is derived from the formulas at the next lower level $L_{k+1}$.
  2. 2.
    Every formula at any level other than the highest (i.e., $k=2,\ldots,K$) is included in exactly one formula at the next higher level $L_{k-1}$.
  3. 3.
    Atomic propositions are used exclusively within the formulas at the lowest level $L_{K}$.
Let $\Phi^{k}$ denote the set of formulas at level $L_{k}$ with $k\in[K]$. For any formula $\phi_{k}^{i}$ at the non-bottom level $k<K$, we slightly bend the notation to use $\phi_{k}^{i}$ to represent the same symbol at the higher level $L_{k-1}$, which we refer to as composite proposition. In this manner, $\phi_{k}^{i}$ represents not only the $i$-th formula at level $L_{k}$, but also the corresponding composite proposition at level $L_{k-1}$. In this work, when $\phi_{k}^{i}$ appears at the right side in a certain formula, we consider it as a composite proposition. On the other hand, when it appears at the left side as a standalone formula, we refer to it as a specification. The direct result of condition 3) is that non-bottom specifications consist solely of composite propositions.
Example 1
continued (Hierarchical sc-LTL) The hierarchical specifications for the task of first picking and placing item $a$ and then fetching items $b$ and $c$ in any order are:
[…]
where there are two levels, $L_{1}$ and $L_{2}$, with $L_{1}$ having one specification and $L_{2}$ having three specifications. Their automatons have a total of 8 states and 18 transitions. The symbol $\phi_{2}^{1}$ is a specification at level $L_{2}$ but is a composite proposition at level $L_{1}$. $\Phi^{1}=\{\phi_{1}^{1}\}$ and $\Phi^{2}=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}$. $\square$
As discussed in Section II-A, local and global forms of task specifications are distinguished based on whether robots receive tasks from individual specifications or from a unified specification. We highlight that the direction distinguishing local from global forms is complementary to the direction distinguishing flat from hierarchical forms. Thus, the hierarchical form can be integrated with both local and global forms. Here, we employ the global form of hierarchical sc-LTL.
Definition IV.2 (Specification Hierarchy Graph)
@@ IV-A Syntax of Hierarchical sc-LTL
The specification hierarchy graph, denoted as ${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$, is a tree where each node represents a specification within the hierarchical sc-LTL, and an edge $(u,v)$ indicates that specification $u$ contains specification $v$ as a composite proposition.
(a) Specification hierarchy graph
(b) NBA of sc-LTL $\Diamond a\wedge\Diamond b$
@@
Figure 1: The specification hierarchy graph corresponds to hierarchical sc-LTL specifications in (2). The NBA corresponds to the sc-LTL specification in Example 2, where non-trivial decomposition states are colored in gray.
@@
Lemma IV.4
Any hierarchical sc-LTL specifications can be turned into a specification hierarchy graph.
@@
continued (Specification hierarchy graph)  The specification hierarchy graph is illustrated in Fig. 1(a). $\square$
IV-B Semantics of Hierarchical sc-LTL
Hierarchical sc-LTL cannot be interpreted by simply substituting composite propositions with their corresponding specifications. For example, replacing composite propositions in Eq. (2) at level $L_{1}$ generates a flat specification:
@@ IV-B Semantics of Hierarchical sc-LTL
continued (Interpretation)  The way of interpretation is to start from the bottom level and progress upwards, verifying if the satisfaction of lower specifications allows the upper ones to move forward. We provide two separate cases. Case (i): If $s_{a}$ now holds true, then $\phi_{2}^{1}$ in (2) transitions to $\Diamond t_{a}$ based on the rules of progression [66] that generate sub-formulas to be satisfied, while $\phi_{1}^{1}$ remains unaffected as none of its composite propositions are fulfilled. If $t_{a}$ subsequently holds true, $\phi_{2}^{1}$ is satisfied. Consequently, $\phi_{1}^{1}$ then simplifies to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. Following this logic, $\phi_{1}^{1}$ will be fulfilled when both $\phi_{2}^{2}$ and $\phi_{2}^{3}$ are satisfied, regardless of the order. Case (ii): Assuming the robot is capable of carrying multiple items. If $s_{b}$ is satisfied, then $\phi_{2}^{2}$ progresses to $\Diamond t_{b}$. Subsequently, as $s_{a}$ and $t_{a}$ sequentially become true, $\phi_{1}^{1}$ transitions to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. At this stage, $\phi_{2}^{2}$ only requires $t_{b}$ to be satisfied. We allow the partial satisfaction of $\phi_{2}^{2}$ provided that it achieves full satisfaction no earlier than $\phi_{2}^{1}$. $\square$
Input: Hierarchical sc-LTL specifications $\Phi$, state-specification plan $\tau$
Output: True or False
1 $\Phi^{acc}=\varnothing$ ;
2 for $i\in\{0,\ldots,h\}$  do
3       for $k\in\{K,\ldots,1\}$  do
4             $\Phi^{acc,k}=\varnothing$ ;
5             for $\phi\in\Phi^{k}\setminus\Phi^{acc}$  do
6                   if $k=K$  then
7                         $S(i)=\left\{s_{r}^{i}\,|\,(s_{r}^{i},\phi)\in\tau^{i}\right\}$ ;
8                         if $S(i)=\varnothing$ then
9                               continue;
10
11                         $\pi^{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\,|\,s_{r}^{i}\in S(i)\right\}$ ;
12
@@ 4             $\Phi^{acc,k}=\varnothing$ ;
14                         $\pi^{i}=\Phi^{acc,k+1}$ ;
15
@@ 14                         $\pi^{i}=\Phi^{acc,k+1}$ ;
16                   ${\mathcal{Q}}_{\mathcal{B}}^{i+1}(\phi)=\left\{q^{\prime}\,|\,(q,\pi^{i},q^{\prime})\in\to_{\mathcal{B}}(\phi),\,\forall q\in{\mathcal{Q}}_{\mathcal{B}}^{i}(\phi),\right.$ $\left.\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi)\right\}$ ;
17                   if ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)\cap{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi)\neq\varnothing$  then
18                         $\Phi^{acc,k}=\Phi^{acc,k}\cup\{\phi\}$ ;
19
20            $\Phi^{acc}=\Phi^{acc}\cup\Phi^{acc,k}\,\cup\,\texttt{Children}(\Phi^{acc,k})$ ;
21
22      if $\phi_{1}^{1}\in\Phi^{acc}$ then
23             return True ;
24
25return False ;
Algorithm 1 Satisfaction check
Given the presence of multiple specifications in hierarchical sc-LTL, a robot could be addressing any specification at a specific state. To clarify the robot’s intention and facilitate the checking of satisfaction, we associate the state $s_{r}$ of a robot $r$ with a leaf specification $\phi_{r}\in\Phi^{K}$. This association forms a state-specification pair $(s_{r},\phi_{r})$, signifying that the robot $r$ is engaged in fulfilling the specification $\phi_{r}$ at state $s_{r}$.
Definition IV.5 (State-Specification Plan)
A state-specification plan with a horizon $h$, represented as $\tau$, is a timed sequence $\tau=\tau^{0}\tau^{1}\tau^{2}\ldots\tau^{h}$. Here, $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ is the collective state-specification pairs of $N$ robots at the $i$-th timestep, where $s_{r}^{i}\in{\mathcal{S}}_{r}$, and $\phi_{r}^{i}\in\Phi^{K}\cup\{\epsilon\}$, with $\epsilon$ indicating the robot’s non-involvement in any leaf specification at that time.
When considering the composite propositions in the context of long-horizon robot tasks, they correspond to sub-tasks. For example, one such task could be first clearing the table and then disposing of the trash, with each sub-task required to be completed exactly once. Another example involves retrieving a beverage from the refrigerator, such as a bottle of apple juice or orange juice, where completing just one sub-task suffices. Given this observation, we propose the following assumption:
Assumption IV.6 (Composite proposition)
For the successful completion of the robotic task, it is adequate that each composite proposition is satisfied no more than once.
Given a robot state-specification plan, the method for determining the satisfaction of hierarchical sc-LTL specifications is shown in Alg. 1. Because the truth of a non-leaf specification depends solely on the specifications at the immediately lower level, the algorithm iteratively assesses specifications in a bottom-up manner. Alg. 1 will stop early and return true if the highest-level specification is met [lines 1-1]. This is because that all specifications belong to co-safe LTL, the subsequent plan becomes irrelevant once the topmost specification is satisfied, eliminating the need for further satisfaction checks. Let $\Phi^{acc}$ represent the set of specifications that have been satisfied up to the current point [line 1]. Additionally, let $\Phi^{acc,k}$ denote the set of specifications at level $L_{k}$ that have just become satisfied at the current time step $i$ [line 1]. According to Assumption IV.6, each composite proposition is adequate to be satisfied no more than once, therefore, we iterate over each specification $\phi\in\Phi^{k}\setminus\Phi^{acc}$ at level $L_{k}$ that remain unsatisfied. Alg. 1 computes the reachable set of automaton states for the next time step, designated as ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)$, to evaluate the satisfaction [lines 1-1]. This assessment varies depending on whether the specification $\phi$ is a leaf one or not. If the specification is a leaf specification, the algorithm aggregates the set of true propositions $\pi^{i}$, which are atomic propositions derived from the states of robots engaged in the specification $\phi$ [lines 1-1]. Otherwise, if the specification is not a leaf, the set of true propositions is identified as the set $\Phi^{acc,k+1}$ – these are the specifications at the immediate lower level that have just been satisfied [line 1]. The set $\Phi^{acc,k}$ gets updated if the reachable automaton states include at least one accepting state of $\phi$ [lines 1]. Following the completion of iterations over level $L_{k}$, the set $\Phi^{acc}$ is then updated [line 1], where $\texttt{Children}(\Phi^{acc,k})$ denotes the set of specifications that are children nodes in the specification hierarchy graph of any specification in $\Phi^{acc,k}$, that is, the set of specifications for which a path exists from any specification in $\Phi^{acc,k}$. There is no need to verify the satisfaction of a specification if its parent specification has already been satisfied. For instance, there is no need to verify the satisfaction of $\phi_{2}^{1}$ for the specification $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2})$ if $\phi_{2}^{2}$ becomes true.

### v4

@@ IV-A Syntax of H-LTLf
There are two levels, $L_{1}$ and $L_{2}$, with $L_{1}$ having one formula and $L_{2}$ having three formulas. The symbol $\phi_{2}^{1}$, appearing on the left side of the equal sign, is a specification at level $L_{2}$ but is a composite proposition at level $L_{1}$ since it is on the right side of the equal sign. $\Phi_{1}=\{\phi_{1}^{1}\}$ and $\Phi_{2}=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}$, $\texttt{Prop}(\phi_{1}^{1})=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}\subseteq\Phi_{2}$ and $\texttt{Prop}(\phi_{2}^{1})=\{s_{a},t_{a}\}\subseteq{\mathcal{A}}{\mathcal{P}}$. This representation reduces the length of each formula, leading to smaller automata and improving interpretability. Their automata have a total of 8 states and 18 transitions. Note that the hierarchical representation is not unique. Depending on the level of granularity in the hierarchy, the same task can be represented as:
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2})$
$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond(s_{a}\wedge\Diamond t_{a}),\quad\phi_{2}^{2}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}$ (3)
$\hskip 28.45274ptL_{3}:\quad$ $\phi_{3}^{1}=\Diamond(s_{b}\wedge\Diamond t_{b}),\quad\,\phi_{3}^{2}=\Diamond(s_{c}\wedge\Diamond t_{c}).\hskip 22.76228pt\square$
Definition IV.2 (Specification hierarchy tree)
@@
The specification hierarchy tree, denoted as ${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$, is a tree where each node represents a specification within the H-LTLf, and an edge $(u,v)$ indicates that specification $u$ contains specification $v$ as a composite proposition. Any H-LTLf specifications can be turned into a specification hierarchy tree.
(a) Specification hierarchy tree
(b) NFA of LTLf $\Diamond a\wedge\Diamond b$
@@
Figure 1: The specification hierarchy tree of H-LTLf specifications in (1) where green nodes represent leaf specifications. The NFA corresponds to the LTLf specification in Example 2, where non-trivial decomposition states are colored in gray.
From a tree perspective, the level $k$ of a specification node $\phi_{k}^{i}$ corresponds to its depth, defined as the number of nodes along the longest path from the root to that node.
@@
Let $\Phi_{\text{leaf}}$ denote the set of leaf specifications. Based on Def. IV.1, leaf specifications consist exclusively of atomic propositions, while non-leaf specifications consist solely of composite propositions.
@@
continued (Specification hierarchy tree)  The specification hierarchy tree is illustrated in Fig. 1(a). $\square$
IV-B Semantics of H-LTLf
H-LTLf cannot be interpreted by simply substituting composite propositions with their corresponding specifications. For example, replacing composite propositions in Eq. (2) at level $L_{1}$ generates a standard specification:
@@ IV-B Semantics of H-LTLf
continued (Interpretation)  The way of interpretation is to start from the bottom level and progress upwards, verifying if the satisfaction of lower specifications allows the upper ones to move forward. We provide two separate cases for specifications in (2). Case (i): If $s_{a}$ now holds true, then $\phi_{2}^{1}$ transitions to $\Diamond t_{a}$ based on the rules of progression [65] that generate sub-formulas to be satisfied, while $\phi_{1}^{1}$ remains unaffected as none of its composite propositions are fulfilled. If $t_{a}$ subsequently holds true, $\phi_{2}^{1}$ is satisfied. Consequently, $\phi_{1}^{1}$ then simplifies to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. Following this logic, $\phi_{1}^{1}$ will be fulfilled when both $\phi_{2}^{2}$ and $\phi_{2}^{3}$ are satisfied, regardless of the order. Case (ii): Assuming the robot is capable of carrying multiple items. If $s_{b}$ is satisfied, then $\phi_{2}^{2}$ progresses to $\Diamond t_{b}$. Subsequently, as $s_{a}$ and $t_{a}$ sequentially become true, $\phi_{1}^{1}$ transitions to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. At this stage, $\phi_{2}^{2}$ only requires $t_{b}$ to be satisfied. We allow the partial satisfaction of $\phi_{2}^{2}$ provided that it achieves full satisfaction no earlier than $\phi_{2}^{1}$. $\square$
Given that H-LTLf includes multiple specifications, we consider multiple systems, particularly transition systems. At any given state, a system may be addressing a particular specification. We associate the state $s_{r}$ of a system ${\mathcal{T}}(r)$ with a leaf specification $\psi_{r}\in\Phi_{\text{leaf}}$. This association establishes a state-specification pair $(s_{r},\psi_{r})$, indicating that at state $s_{r}$, the system ${\mathcal{T}}(r)$ is engaged in satisfying the leaf specification $\psi_{r}$.
Definition IV.4 (State-Specification Sequence)
A state-specification sequence with a horizon $h$, represented as $\tau$, is a timed sequence $\tau=\tau_{0}\tau_{1}\tau_{2}\ldots\tau_{h}$. Here, $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ is the collective state-specification pairs of $N$ systems at the $i$-th timestep, where $s_{r}^{i}\in{\mathcal{S}}_{r}$, and $\psi_{r}^{i}\in\Phi_{\text{leaf}}\cup\{\epsilon\}$, with $\epsilon$ indicating the system’s non-involvement in any leaf specification at that time.
Given the state-specification sequence $\tau$ and a leaf specification $\phi$, we generate a word as $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$, where $\sigma_{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\ |\ (s_{r}^{i},\psi_{r}^{i})\in\tau_{i}\;\text{and}\;\psi_{r}^{i}=\phi\right\}$, represents the collective observations generated by the systems engaging in the leaf specification $\phi$ at instant $i$. Furthermore, given a leaf (non-leaf) specification $\phi$ and an input word $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ with $\sigma_{i}\subseteq\texttt{Prop}(\phi)$, consisting of its atomic (composite) propositions, we construct an output word of $\phi$, denoted by $\sigma^{\prime}=\sigma^{\prime}_{0},\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n}$, where $\sigma^{\prime}_{i}$ is either $\varnothing$ or $\{\phi\}$, tracking the satisfaction of $\phi$. Let $\sigma_{i,j}=\sigma_{i},\ldots,\sigma_{j}$ denote the segment of the input word between indices $i$ and $j$.
Definition IV.5 (Output word)
Given a specification $\phi$ and an input word $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ with $\sigma_{i}\subseteq\texttt{Prop}(\phi)$, the output word $\sigma^{\prime}=\sigma^{\prime}_{0},\sigma_{1}^{\prime},\ldots,\sigma^{\prime}_{n}$ of $\phi$ is defined as:
  • By default, $\sigma^{\prime}_{-1}=\{\phi\}$.
  • $\sigma^{\prime}_{j}=\{\phi\}$ if $\sigma_{i+1,j},i+1\models\phi$, where $i$ (with $i<j$) is the most recent instant when $\phi$ was satisfied, i.e., $i=max\{k\,|\,\sigma^{\prime}_{k}=\{\phi\},k=-1,0,\ldots,j-1\}$.
The satisfaction relation $\models$ follows the same definition as introduced in Section III, treating composite propositions as atomic propositions for non-leaf specifications. The progress of $\phi$ resets after it is satisfied at instant $i$, and it is considered satisfied again at instant $j$ if the input word segment $\sigma_{i+1,j}$ satisfies $\phi$. In the context of robots, this implies that the duration of task $\phi$ extends from $i+1$ to $j$, with the completion time at instant $j$. The task’s completion status is not persistent, meaning that once completed, it must start from scratch if repeated. The output words of specifications at level $L_{k+1}$ serve as the input words and produce output words of specifications at level $L_{k}$. Thus, starting from the words of leaf specifications generated by the state-specification sequence $\tau$ and proceeding in a bottom-up manner, we ultimately obtain an output word of the root specification $\phi_{1}^{1}$. The whole H-LTLf is satisfied by a state-specification sequence $\tau$ if the output word of $\phi_{1}^{1}$ contains $\phi_{1}^{1}$. Algorithmically, a post-order traversal can be used to traverse the specification hierarchy tree to compute the output word of $\phi_{1}^{1}$, as outlined in Alg. 1.
Definition IV.6 (Semantics)
[…]
Input: state-specification sequence $\tau$, LTLf  specification $\phi$, H-LTLf specifications $\Phi$
Output: output word $\sigma^{\prime}$ of specification $\phi$
  $\triangleright$ Get input word $\sigma$ [lines 1-1]
  $\triangleright$ Leaf specification
1 if $\phi\in\Phi_{\text{leaf}}$ then
2       $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ where $\sigma_{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\ |\ (s_{r}^{i},\psi_{r}^{i})\in\tau_{i}\;\text{and}\;\psi_{r}^{i}=\phi\right\}$;
3
  $\triangleright$ Non-leaf specification
@@
5       for $\phi^{\prime}\in\texttt{Prop}(\phi)$ do
6             $\sigma(\phi^{\prime})=\texttt{GenerateOutputWord}(\tau,\Phi,\phi^{\prime})$ ;
7
        $\triangleright$ Merge output words at the lower level as the input word of $\phi$
@@ 6             $\sigma(\phi^{\prime})=\texttt{GenerateOutputWord}(\tau,\Phi,\phi^{\prime})$ ;
8       $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ where $\sigma_{i}=\left\{\sigma_{i}(\phi^{\prime})\;|\;\forall\phi^{\prime}\in\texttt{Prop}(\phi)\right\}$ ;
9
  $\triangleright$ Get output word $\sigma^{\prime}$ [lines 1-1]
$i=-1$ ;
  $\triangleright$ Latest time when $\phi$ is true
10 for $j\in\{0,1,\ldots,n\}$ do
11       if $\sigma_{i+1,j},i+1\models\phi$ then
12             $\sigma^{\prime}_{j}=\{\phi\}$;
13             $i=j$;
14
15      else
16             $\sigma^{\prime}_{j}=\varnothing$ ;
17
18return $\sigma^{\prime}$;
Algorithm 1 GenerateOutputWord($\tau,\Phi,\phi$)
$t:$00111122223333444455556666777788889999obsv:$\{s_{a}\}$$\{t_{a}\}$$\{s_{b}\}$$\{t_{b}\}$$\{s_{c}\}$$\{t_{c}\}$ϕ21::superscriptsubscriptitalic-$\phi_{2}^{1}:$$\{\phi_{2}^{1}\}$ϕ22::superscriptsubscriptitalic-$\phi_{2}^{2}:$$\{\phi_{2}^{2}\}$ϕ23::superscriptsubscriptitalic-$\phi_{2}^{3}:$$\{\phi_{2}^{3}\}$ϕ11::superscriptsubscriptitalic-$\phi_{1}^{1}:$$\{\phi_{1}^{1}\}$
Figure 2: Each row illustrates either the observation or the output word of a specification. The filled nodes indicate scenarios where an atomic or composite proposition is satisfied. The observations (obsv) generate the input word for leaf specifications $\phi_{2}^{1}\sim\phi_{2}^{3}$, while the combination of output words for specifications $\phi_{2}^{1}\sim\phi_{2}^{3}$ serve as the input word for the non-leaf specification $\phi_{1}^{1}$. The dashed arrows illustrate the correspondence between specific inputs that lead to the satisfaction of a given specification.
Example 1
continued (Semantics)  The output words for the specifications in (2) are illustrated in Fig. 2. $\square$